# Theoretical framework for lattice QCD computations of $B\to K\ell^{+}\ell^{-}$ and $\bar{B}_{s}\to\ell^{+}\ell^{-}\gamma$ decays rates, including contributions from “Charming Penguins”

### v1

@@ VIII.3 Comparison with a simple model for the spectral density and the extrapolation to $\varepsilon=0$
We find that in order to match the model predictions to the lattice data below threshold ($m<m_{J/\psi}+m_{\eta_{ss^{\prime}}}$) for the $O_{1}^{(c)}$ contribution an additional multiplicative factor of about $2$ needs to be introduced. This is reasonable, as the typical magnitude of the missing renormalization is expected to be an $O(1)$ effect. In the figures the model predictions have been multiplied by this factor. Overall, as the figure shows, we do observe a fairly good agreement between the model predictions and the lattice results for the contribution from the operator $O_{1}^{(c)}$. The analogous comparison for the $O_{2}^{(c)}$ contribution to $B_{s}\to\eta_{ss^{\prime}}\ell^{+}\ell^{-}$ is shown in Fig. 17 for $\alpha=1.5$ and $0.8$.

### v2

@@ VIII.3 Comparison with a simple model for the spectral density and the extrapolation to $\varepsilon=0$
We find that in order to match the model predictions to the lattice data below threshold ($m<m_{J/\psi}+m_{\eta_{ss^{\prime}}}$) for the $O_{1}^{(c)}$ contribution an additional multiplicative factor $2.3$ needs to be introduced. This is reasonable, as the typical magnitude of the missing renormalization is expected to be an $O(1)$ effect. In the figures the model predictions have been multiplied by this factor. Overall, as the figure shows, we do observe a fairly good agreement between the model predictions and the lattice results for the contribution from the operator $O_{1}^{(c)}$. The analogous comparison for the $O_{2}^{(c)}$ contribution to $B_{s}\to\eta_{ss^{\prime}}\ell^{+}\ell^{-}$ is shown in Fig. 17 for $\alpha=1.5$ and $0.8$.